how many Hydronium Ions in 1.5L water?
Answer #1

[H3O+] in water = 1.0*10^-7 M

So,

Number of mol of H3O+ = M * volume

= 1.0*10^-7 M * 1.5 L

= 1.5*10^-7 mol

Number of H3O+ = Number of mol of H3O+ * Avogadro’s number

= 1.5*10^-7 * 6.022*10^23

= 9.03*10^16

Answer: 9.0*10^16